OpenStudy (magbak):

I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!!!!!!! PLEASE HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Find the 6th partial sum of 78,432 11,204 9,555 2,387 @whpalmer4

OpenStudy (anonymous):

6th partial sum of ....

OpenStudy (magbak):

\[\sum_{i=1}^{\infty}7(4)^i-1\]

OpenStudy (magbak):

Bam that is it

OpenStudy (whpalmer4):

You're in good hands, and I hear a pizza calling my name...

OpenStudy (magbak):

Yes I am the pizza.

OpenStudy (anonymous):

6th partial sum is \[\sum_{i=1}^67(4)^i-1\]

OpenStudy (anonymous):

which you can compute by grinding it out or writing \[\sum_{i=1}^67(4)^i-1=7\sum_{i=1}^6(4)^i-\sum_{i1}^61\]

OpenStudy (anonymous):

then \[\sum_{i=1}^64^i=\frac{4^7-4}{4-1}=5460\] making your answer \[7\times 5460-7\]

OpenStudy (whpalmer4):

that's not among the answer choices, @magbak, are you sure you've given us the right sum to evaluate?

OpenStudy (whpalmer4):

that should be \[7*5460-6\]FWIW

OpenStudy (anonymous):

oops

OpenStudy (magbak):

Yes

OpenStudy (magbak):

No one sec

OpenStudy (magbak):

\[\sum_{i=1}^{\infty}7(4)^(x-1)\]

OpenStudy (magbak):

@satellite73

OpenStudy (primeralph):

"Yes I am the pizza"?? Sounded gay.

OpenStudy (anonymous):

lost me now i don't understand what you wrote is it \[\sum_{i=1}^67\times 4^{i-1}\]

OpenStudy (magbak):

Yes

OpenStudy (magbak):

Oh no he said the pizza was calling and I am calling for his help no homo bro.

OpenStudy (anonymous):

ok then pull out the 4, get out a calculator and add \[7(1+4+4^2+4^3+4^4+4^5)\]

OpenStudy (magbak):

Ok thanks one more question and I am done ok.